OpenStudy (anonymous):

what is a simpler form of the expression? (2n + 5n + 3)(4n - 5)

OpenStudy (asnaseer):

you can sum the 2n and 5n to simplify

OpenStudy (anonymous):

(2n + 5n + 3)(4n - 5)=(7n+ 3)(4n - 5) this is what asnaseer meant by sum... now just multiply them example: \[\LARGE (a+b)(c-d)=ac-ad+bc-bd\] in your case you have: \[ \LARGE (7n+ 3)(4n - 5)=... \]